OpenStudy (angela210793):

2^(a-b)=6 2^(a+b)=12 find b=?

OpenStudy (anonymous):

well just try logarithm?

OpenStudy (anonymous):

log2,6=a-b log 2,12=a+b

OpenStudy (anonymous):

wait i take my notebook and my pencil ^^

OpenStudy (angela210793):

Can we do it like this: 2^(a-b)=6 -----------=2^(a-b-a-b)=1/2 2^(a+b)=12 2^-2b=1/2 2^-2b=2^-1 -2b=-1 b=1/2

OpenStudy (angela210793):

I don't think we need log....

OpenStudy (anonymous):

m2 just its a way ...

OpenStudy (amistre64):

you can do that since the bottom w know dont equal 0

OpenStudy (angela210793):

kk thnx ^_^ @Amistre can u pls look 1 of my q?

OpenStudy (amistre64):

which one?

OpenStudy (anonymous):

its simple 2^a - 2^b = 6 2^a + 2^b=12

OpenStudy (angela210793):

this 1: What should m be so that : (m+3)x^2-5mx+4m+1>0 for every x...

OpenStudy (anonymous):

and now its 2 x 2^a = 18 ^^

OpenStudy (angela210793):

Korcan 2^(a-b)=2^a/2^b

OpenStudy (amistre64):

(m+3)x^2 +(-5m) x + (4m +1) > 0 its a quad that needs to be accounted for a b and c

OpenStudy (amistre64):

\[x = 0; when;x=\frac{5m \pm \sqrt{25m^2 -4(m+3)(4m+1)}}{2(m+3)}\]

OpenStudy (amistre64):

we can narrow down the results by working the discriminate

OpenStudy (angela210793):

yea i found this 2..but i can't go any further...O.o

OpenStudy (anonymous):

u see u are better then me...

OpenStudy (angela210793):

right..:P same story :P

OpenStudy (amistre64):

25m^2 -4m^2 -16m -12 >= 0

OpenStudy (amistre64):

21m^2 -16m -12 >= 0 ; quad form that to see your options

OpenStudy (angela210793):

tht's it? Thnx ^_^ :)

OpenStudy (anonymous):

well angela can you help me on my question?

OpenStudy (angela210793):

hmmmm..idk..let me see

OpenStudy (amistre64):

the roots to that are: m= \[\frac{2}{21}(4\pm \sqrt{79})\]

OpenStudy (amistre64):

...... i missques the discriminate when I did that :/

OpenStudy (angela210793):

Sorry but i need some english lessons too..wht's missques? O.o